I don't get this one at all. Solve using the zero product property. 2x^4-16x=0
When I tried to solve it, I could get the zero and two answers but they also showed an answer with a number +- square root and I must be doing something wrong because I am not coming up with it. Any help in the right direction would be greatly appreciated. Thanks.

Answer by stanbon(75887)

2x^4-16x=0
---
Factor:
2x(x^3-8) = 0
x(x^3-8) = 0
x(x-2)(x^2+2x+4) = 0
------------------------------
x = 0, or x = 2 or x = the solutions for the quadratic:
x = [-2+-sqrt(4-4*4)]/2
x = [-2+-sqrt(-12)]/2
x = [-2+-2isqrt(3)]/2
x = [-1+-isqrt(3)]
--------------------------
Cheers,
Stan H.